OpenStudy (anonymous):

Fool's problem of the day (Run down memory lane) After the successful completion of the today's problem, I thought of posting five of the old problems that haven't been (correctly) answered, 1) Prove that \( \tan(20^\circ) \cdot \tan(40^\circ) \cdot \tan(80^\circ)= \tan(60^\circ) \) [Solved by @Mani_Jha] 2) How many eight digit numbers formed using the digit \( \{7,8,9\} \) has exactly three \( 7 \)'s ? [Solved by @PaxPolaris] 3) what is the remainder when we divide \( \large 9^{{100}^{{101}^{102}}}-5^{81^{82^{83}}} \) by \(13\). [Solved by @KingGeorge] (Note: There is a far more simpler yet analytic way of solving this!) [EDIT: The simpler way is also evinced by @kingGeorge] 4) If \( n \) and \( m \) are natural numbers such that \( 1 + 2 + 3 + … + n = m^2 \), Find the sum of digits of the greatest such \( n \) , smaller than \( 10^3 \). [PS:This has a very simply non-programming solution] 5) Prove that \( \sum \limits_{k=0}^{n-r} = \binom{n}{k} \times \binom{n}{r+k} = \binom{2n}{n-r} \) [Excellent hint by @satellite73 and solved by @Mani_Jha] PS: It's safe to assume that none of these are very easy. Good luck!

OpenStudy (anonymous):

3) -11 or 2

OpenStudy (anonymous):

FoolForMath, Makes the site more Fun:)

OpenStudy (anonymous):

@cinar: I am sorry, it's not.

OpenStudy (anonymous):

Thanks @Open2study

OpenStudy (turingtest):

Will I ever solve one of your problems FFM?

OpenStudy (experimentx):

Let's keep trying ...!

OpenStudy (anonymous):

\[9^3=1(\mod13)\] \[5^4=1(\mod13)\]

OpenStudy (paxpolaris):

2) number of choosing a 5 digit number with all being 8 or 9 is \(\large 2^5\) ...times number of way to choose 3 positions for the 7s.....\(\large (_8C_3)\) \[\large2^5\left(\begin{matrix}8 \\ 3\end{matrix}\right)=32\left( 8\cdot7 \right)=\huge1792\]

OpenStudy (inkyvoyd):

Oh no! right after I post my cheap rip-off geometry problem!

OpenStudy (anonymous):

@TuringTest: You can! I am sure you will :)

OpenStudy (anonymous):

4) 8

OpenStudy (anonymous):

4) \[\sqrt{\frac{n(n+1)}{2}}=m=\mathbb{N}\]

OpenStudy (paxpolaris):

4) n=49 is too big

OpenStudy (kinggeorge):

2) To find this, we first need to calculate\[\large 100^{101^{102}} \mod \phi(13)\equiv 100^{101^{102}} \mod 12\]For this we need to calculate\[101^{102} \equiv 1^2\equiv 1\mod 4\]Since \(\phi(12)=4\).That means that \[\large 100^{101^{102}} \mod 12\equiv 100^1 \equiv 4 \mod 12\]Finally, we can find\[\Large 9^{100^{101^{102}}} \equiv 9^4 \mod 13\equiv 9\]

OpenStudy (kinggeorge):

Now we have to do the second part of that problem.

OpenStudy (kinggeorge):

To find \[\Large 5^{81^{82^{83}}} \mod 13\]We need to find \[\large {81^{82^{83}}} \mod 12\]So we need to find \[\large 82^{83} \equiv 2^3 \equiv 0\mod 4\]This means that \[81^{82^{83}} \equiv 9^0 \equiv 1 \mod 12\]Finally, \[\Large 5^{81^{82^{83}}} \equiv 5^1\equiv 5 \mod 13\]

OpenStudy (anonymous):

4) could be 9 1 and 8

OpenStudy (kinggeorge):

Now that we have both parts, it's simple to see that \[\large 9-5 \equiv 4\mod 13\] So the answer to 3) should be 4.

OpenStudy (paxpolaris):

4) @Arnab09 is right: n=8, m=6 sum is 36

OpenStudy (anonymous):

@paxpolaris how about n=1 m=1

OpenStudy (paxpolaris):

largest sum below 1000

OpenStudy (anonymous):

I think the best way to solve #4 is programing

OpenStudy (paxpolaris):

4) 45 not perfect square???

OpenStudy (anonymous):

I tried from n=1 to n=50 bit it takes too long to other numbers..

OpenStudy (experimentx):

from 1 to 2000 we have 1,2,9,1,2,9,50,289,1682

OpenStudy (experimentx):

I think answer will be 9 since the sum is less than 100, next occurance will be at 50 where sum > 1000

OpenStudy (paxpolaris):

@experimentX i think u need to subtract 1 on all nos.

OpenStudy (anonymous):

@expermentx you put n=1000 in your program right?

OpenStudy (experimentx):

i put 2000

OpenStudy (experimentx):

but why did i get 2 i don't know ..

OpenStudy (anonymous):

I am not good at programing :(

OpenStudy (experimentx):

@cinar change it into sine and cosines ... i remember it doing.

OpenStudy (kinggeorge):

Second method of doing 3: We know that \(9^3 \equiv 1 \mod 13\) and that \(5^4\equiv 1\mod13\)So we only need to calculate \[\large 100^{101^{102}} \equiv 1^{2^0} \equiv 1\mod 3 \]and\[81^{82^{83}} \equiv 1^{2^{3}} \equiv 1 \mod 4\]Thus the problem is reduced to \[9^1-5^1 \equiv 9-5\equiv 4\mod 13\]So we're done.

OpenStudy (kinggeorge):

@FoolForMath Was my second solution the simpler way of doing it?

OpenStudy (anonymous):

@KingGeorge: You are right.

OpenStudy (anonymous):

last one maybe a verbal description after writing \(\binom{2n}{n-r}=\binom{2n}{n+r}\)

OpenStudy (anonymous):

i have made some progress with 1

OpenStudy (anonymous):

i have shown cos20cos40cos80 = 1/8

OpenStudy (experimentx):

sin20sin40sin80 = sqrt(3)/8

OpenStudy (anonymous):

has someone already completed 1?

OpenStudy (experimentx):

most likely, for problem 4) either (n and n+1/2) must be a perfect square or ( n/2 and n+1) must be a perfect square. the possibilities are 1 4 9 16 25 36 49 for n, n/2, n+1, n+1/2 are these ... it would simplify our job a lot easier than programming. let n=1, then (n+1)/2 = 1 => so n= 1, sum = 1 sum^2 = 1 let n/2=1, then (n+1) => 3 => nps let n+1=1, n/2=0 => nps let (n+1)/2=1, n=1, n=1, sum^2 = 1 let n=4, then (n+1)/2 => 5/2 => nps let n/2 = 4, then (n+1) => 9 => sum = 36 let n+1 = 4, then n=5 => nps let (n+1)/2 = 4, then n=7 => nps let n=9, then (n+1)/2=5 => nps let n/2=9, then (n+1) = 19 => nps let (n+1)/2 = 9, then n=17 => nps let n+1 = 9, n/2=4 => sum = 36 let n=16, then (n+1)/2=17/2 => nps let n/2=16, then (n+1) = 33 => nps let (n+1)/2 = 16, then n=31 => nps let n+1 = 16, then n/2=15/2 => nps let n=25, then (n+1)/2=13 => nps let n/2=25, then (n+1) = 51 => nps let (n+1)/2 = 25, then n=49 => sum is greater than 1000 let n+1 = 25, then n/2=12 => nps similarly we can check for 36 the answer would be 8

OpenStudy (anonymous):

i am now trying to show that

OpenStudy (anonymous):

how did you do it? im stuck :/

OpenStudy (experimentx):

@eigenschmeigen i don't think anyone has ... i saw the proof for sin20sin40sin60 somewhere ... i won't be posting answer

OpenStudy (anonymous):

proof for the cos: \[\sin2\theta = 2\sin\theta \cos\theta\] \[\frac{\sin2\theta}{ 2\sin\theta}= \cos\theta\] \[cos\theta cos2\theta cos 4\theta = \frac{sin2\theta sin4\theta sin8\theta}{8sin\theta sin2\theta sin4\theta} \] let\[\theta = 20\] \[\sin{160} = \sin20\] so we get: \[cos20cos40cos80= \frac{1}{8} \]

OpenStudy (anonymous):

Problem 4) solution is attached.

OpenStudy (anonymous):

so aswer for #4 is 346

OpenStudy (dumbcow):

i can prove #1 real quick with my calculator :)

OpenStudy (callisto):

@cinar according to @robtobey , n=288 sum = (1+288)*288/2 @dumbcow Good 'method', I've never thought of it though :(

OpenStudy (inkyvoyd):

Actually, it probably isn't.

OpenStudy (inkyvoyd):

http://en.wikipedia.org/wiki/Square_triangular_number FOR NUMBER 4.

OpenStudy (inkyvoyd):

@FoolForMath ?

OpenStudy (anonymous):

i dont think the idea was to google the problems :P

OpenStudy (anonymous):

and @dumbcow that wouldn't be proving it, as the calculator would have inaccuracy

OpenStudy (experimentx):

i agree .. with eigenschmeigen

OpenStudy (anonymous):

i am determined to prove the sine part today

OpenStudy (anonymous):

it will happen

OpenStudy (anonymous):

i will make it so

OpenStudy (experimentx):

best of luck ... i saw it on internet ... beware of google.

OpenStudy (dumbcow):

i was only half-kidding, yes i realize that is not a mathematical proof

OpenStudy (mani_jha):

\[\tan(60+A)\times tanA \times \tan(60-A)=\tan3A\] This identity works for the first one. Taking A =20, we write 80=60+20 and 40=60-20 So the result is tan (3*20)=tan60

OpenStudy (anonymous):

@Mani_Jha: Congratz! :) Do you know how the prove that identity itself?

OpenStudy (mani_jha):

I can prove that identity for the sine and cosine functions.(using the sin(A+-B) formulae) Then I can divide to obtain that identity.

OpenStudy (mani_jha):

For the last problem, the first term represents the sum of binomial coefficients of : \[(1+x)^{n}=\left(\begin{matrix}n \\ 0\end{matrix}\right)+\left(\begin{matrix}n \\ 1\end{matrix}\right)\]........ The second term represents the sum of the binomial coefficients of a part of the same series: \[(1+x)^{n}=....\left(\begin{matrix}n \\ n-r\end{matrix}\right)x ^{n-r}+\left(\begin{matrix}n \\n-r-1\end{matrix}\right)\]........... When we multiply both all we have to do is to collect the coefficients of x^(n-r) in the multiplication of the series: That will be: \[nC0\times nCn-r+nC1 timesnCn-r-1...........\] which leads to the Left hand side(the sum of the subscripts should always be n-r) The problem is simply collecting the coeffecient of x^n-r in the expansion of: \[(1+x)^{2n}\]

OpenStudy (inkyvoyd):

I give up easily, then I just use google. Most of these problems are beyond me anyways. I haven't actually taken any formal math beyond algebra two.